OpenStudy (anonymous):

simplify (3x+3/x-2)-(x-4/x-2)

OpenStudy (anonymous):



OpenStudy (anonymous):

mayebe u can say thanks :)

OpenStudy (anonymous):

i vote best answer and give a medle im sorry i thought that was a good thank you

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

thats not one of my options tho

OpenStudy (anonymous):

put all your options

OpenStudy (campbell_st):

slight problem with the solution... both fractions have the same denominator so its \[\frac{3x + 3 - x + 4}{x -2} = \frac{2x + 7}{x-2}\]